40ml of x M kMnO4 solution is required to react completely with 200ml of 0.02M oxalic acid solution in acidic medium. The value of x is


1
Expert's answer
2021-07-05T07:46:29-0400

"Molarity= moles\/litre"


40ml = 40g

Moles= 40/158= 0.253moles


X= 0.253/1000

X= 2.5× 10-4 M